OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{(-2)^{n}}{n^{n}}\] = \[\sum_{n=1}^{\infty} \left( \frac{-2}{n}\right)^{n}\] = Root test \[\frac{-2}{n} \] Doesn't really work :( The Ratio test, on the other hand, leaves me with \[\frac{(-2)(n^{n})}{(n+1)^{n+1}}\]

OpenStudy (turingtest):

continue with the ratio test I say

OpenStudy (turingtest):

\[\frac{(-2)(n^{n})}{(n+1)^{n+1}}=-2{n^n\over n^{n+1}+O(n)^k}\]where\[k\le n\]

OpenStudy (turingtest):

so divide top and bottom by \(n^n\)

OpenStudy (anonymous):

\[ (-2) \left( \frac{n}{n+1} \right)^{n}\]

OpenStudy (anonymous):

sorry just had to write that thought down...anyhow

OpenStudy (anonymous):

becomes 1/n

OpenStudy (anonymous):

so zero

OpenStudy (anonymous):

Agreed?

OpenStudy (turingtest):

yes, sorry, my power went out :/